Grade 12th passIIT JEE Entrance Exam

A passenger and a goods train are headed in the same direction on parallel tracks the passenger train is 240 metre long and has a constant velocity of 72 km per hour beginning from the time the engine of the passenger train approaches The Last wagon of the goods train it takes 25 second to be in the level with the engine of the goods train it to 30 seconds more to completely overtake the goods train determine the length and the speed of the goods train

Bhavya Srivastava
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
 
Let the length and velocity of goods train to be ‘y’ m  and ’x’ m/s respectively.
Now, velocity of passenger train = 72km/h = 72* 5/18 = 20 m/s.
Also, given length of passenger train = 240 m.
Relative velocity of passenger train with respect to goods train = (20 – x) m/s.
Distance covered by passenger train with respect to goods train in 25 s is given by:
y = (20 – x)* 25------(1)
Also, distance covered by passenger train with respect to goods train to overtake it in 30 s is given by :
240 = (20 – x)* 30------(2)
Now, dividing equation (1) by (2) on both the sides simultaneously, we get :
y/240 = 25/30
On solving the above equation we got :
y = 200 m
Hence, the length of goods train is 200 m.